OpenStudy (anonymous):

Please help!! The largest Ferris Wheel in the world is the London Eye in England. The height (in meters) of a rider on the London Eye after t minutes can be described by the function h(t) = 65 sin[12(t − 7.5)] + 70. (15) A) At what time(s) will the rider be at a height of 100 m? B) How long does it take for the Ferris wheel to go through one rotation?

OpenStudy (aum):

h(t) = 65 sin[12(t − 7.5)] + 70 "At what time(s) will the rider be at a height of 100 m" Set h(t) = 100 and solve for t: 100 = 65 sin[12(t − 7.5)] + 70

OpenStudy (princeharryyy):

u r back @Katieholmes

OpenStudy (anonymous):

Yes im back! @princeharryyy

OpenStudy (princeharryyy):

gud @Katieholmes i think u have got ur assignments back

OpenStudy (anonymous):

which assignment?

OpenStudy (princeharryyy):

watever myt be given to u dat I don't kno

OpenStudy (anonymous):

no, no assignments just studying for the exam

OpenStudy (princeharryyy):

ohk from which date do they start

OpenStudy (anonymous):

september 1st

OpenStudy (princeharryyy):

ohk ... i hate them ...

OpenStudy (princeharryyy):

they make us paranoid

OpenStudy (aum):

For b): If the function is Asin(Bx - C) + D, the period is (2*pi) / B. How long does it take for the Ferris wheel to go through one rotation? They are asking for the period. 65 sin[12(t − 7.5)] + 70. Here B = 12. Therefore, the period = (2*pi)/12 = pi/6 = 0.5236 minute.